Find the y-intercept and the x-intercept for the slope: y=-2x^2+4x+2
Answer by ewatrrr(24785)


Hi
y = -2x^2+4x+2 |Note use of ^ for squared term
y-intercept (when x = 0) Pt(0,2)
y= -2(x-1)^2 + 4 |completing the square Vertex Pt(1,4)
x-intercept (when y = 0)
-2(x-1)^2 + 4 =0
(x-1)^2 = 2
x = 1 ± sqrt(2) Pt(2.414,0) and Pt(-.414,0) the x-intercepts